OpenStudy (anonymous):

I'm already stuck on this problem What is the 7th term of the geometric sequence where a1 = 625 and a2 = -125? First term: 625

hartnn (hartnn):

common ratio = a2/a1 =.. ?

OpenStudy (anonymous):

-125/625

OpenStudy (anonymous):

= -5

hartnn (hartnn):

thats your 'r' a1 =625 we want 7th term, so n=7

hartnn (hartnn):

-5, sure ? check again ....

OpenStudy (anonymous):

dont tell m the answer yet!!!!!!

OpenStudy (anonymous):

me*

OpenStudy (anonymous):

-0.2

OpenStudy (anonymous):

R = -0.2

hartnn (hartnn):

r= -0.2 or -1/5 is correct. now use the formula

OpenStudy (anonymous):

ok!! no answers yet please

OpenStudy (anonymous):

0.04000000000000002

OpenStudy (anonymous):

in other words: (or numbers) 0.04?

OpenStudy (anonymous):

The formula for the nth term in a Geometric sqeuence: \[a_n=a_1r^{n-1}\] Plugging in \(a_1=625\) we have \[a_n=625r^{n-1}\] For the second \((n=2)\) term of the sequence we have \(a_2=-125\). Lets plug \(a_2\) \[a_2=625r^{2-1} \rightarrow-125=625r^1\] We now have \(r\)!

hartnn (hartnn):

0.04 is correct :) no point in writing so many zero's :P

OpenStudy (anonymous):

oh my gosh i did it all by myself!!!! thank you for your help....you taught me well!!!!

hartnn (hartnn):

*applauses*

OpenStudy (anonymous):

thank you thank you!!!!!!

hartnn (hartnn):

welcome welcome!!!!!!